What Effort Will Be Required to Lift a Load of 500 N by a Single Movable Pulley? Hint: Mechanical Advantage of a Single Movable Pulley is Two



1
Expert's answer
2021-03-29T08:57:37-0400

For the single movable pulley

"\frac{F}{F_1}=2\implies"

"F_1=\frac{F}{2}=\frac{500}{2}=250~N."